A 500 mm-diameter concrete pipe is laid on a slope of 1m per 500m and is required to carry water at the rate of 0.04m^3/s.Determine the normal depth of flow.Use roughness coefficient n=0.013


1
Expert's answer
2021-04-14T07:46:19-0400

Q= flow rate = 0.04 "m^3\/s"

R= hydraulic radius

s= slope ("\frac{1}{500})"

n= manning's coefficient of roughness= 0.013

"R= \frac{A}{p} = \frac{weted area}{weted perimeter} = \frac{0.5h^2}{3.412h}"

"Q=VA"


"V= \frac{1}{(n)} (R^ \frac{2}{3}S ^ \frac{1}{2})"

"Q= \frac{1}{(n)} (\frac{0.5h^2}{3.412h})^\frac{2}{3} \times (\frac{1}{500})^\frac{1}{2} \times 0.5h^2"

"0.04= \frac{1}{(0.013)} (\frac{0.5h^2}{3.412h})^\frac{2}{3} \times (\frac{1}{500})^\frac{1}{2} \times 0.5h^2"


solving for h,

h , normal depth is equal to, h= 0.1854 m